The width of a rectangular lot is 75% of its length. If the perimeter is 700 meters, then what are the length and width?

Answer by mananth(16949)

The width of a rectangular lot is 75% of its length. If the perimeter is 700 meters, then what are the length and width?
...
let length =x
width = 3x/4
perimeter = 2*(L+W)
2*(x+3x/4)=700
/2
7x/4 = 350
multiply by 4
7x=350*4
7x=1400
x= 200m the length
width = 3/4*200
=150m
...
m.ananth@hotmail.ca